OpenStudy (skullcandy229):

If A = -3n + 2 and B = 5n - 7, what is the value of A - B, in simplest form? I will give you a medal

OpenStudy (anonymous):

A - B = (-3n + 2) - (5n - 7) = (-3n - 5n) + (2 - (-7)) = (-8n) + (9) = 9 - 8n